A solenoid of length 2.2 m has a diameter of 0.9 m. The magnetic field at its centre is 0.4 tesla. Estimate the energy stored in the field of this coil. 


1
Expert's answer
2021-01-12T12:13:25-0500

The energy stored in the solenoid is

"U=\frac{B^2 Al}{2 \mu_0}"

"=\frac{B^2 \pi r^2l}{2 \mu_0}" "=\frac{(0.4)^2 \pi (0.45)^2(2.2)}{2 (4 \pi \times10^{-7})}" "=2.23 \times 10^5 J"